OpenStudy (anonymous):

Factor completely. Show all work. If polynomial is prime state this. 6x^2+21x+15

OpenStudy (anonymous):

(2x + 5)(3x+3)

OpenStudy (anonymous):

Can you show work on it, so I can understand it better?

OpenStudy (anonymous):

its just factoring, using the factors of 6 and 15 when you try to foil them they have to equal the middle number of 21

OpenStudy (anonymous):

oh ok, thank you! I knew it was easier than it looked.

OpenStudy (anonymous):

I foiled it out and this is my work, is it right? 2x*3x =6x^2 5*3=15 (2x+5)(3x+3) F=6x^2 O=5x I=15x L=15 6x^2+20x+15 How come the middle number is not 21? or is this right?

OpenStudy (anonymous):

I got it, thank you, I did the O wrong, it should be 6x! Thanks.